OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above?

OpenStudy (anonymous):

A. y = - x + 30B. y = 3x + 6C. y = - x + 6D. y = x + 30

OpenStudy (anonymous):

?

OpenStudy (anonymous):

B